# On uniqueness of radial potentials for given Dirichlet spectra with distinct angular momenta

### v1

@@ Appendix A A Bessel basis of $L^{2}(0,1)$
Substituting into (A.16) and (A.17), we deduce that
$b\sim 25.8125\,K,\qquad c\sim-128.293\,K.$
Plugging these values into (A.19), this implies that $K=0$ and hence $b=c=0$. Thus, $f(t)=a$. At last, since $f$ is orthogonal to the constants, $a=0$, which completes the proof of the lemma. ∎
From the preceding lemma, we deduce the following result:
Theorem A.2.
The family
$\left(1,\ \bigl(\Phi_{0}(n\pi\cdot)\bigr)_{n\geq 1},\ \bigl(\Phi_{2}((n+1)\pi\cdot)\bigr)_{n\geq 1}\right)$
forms a basis of $L^{2}(0,1)$.
Proof.
The argument follows the proof of Corollary 5.2 in [23]. Assume that there exist coefficients $\{c_{n}\}_{n\geq 1}$, $\{d_{n}\}_{n\geq 1}$ and $d_{0}\in\mathbb{R}$ such that
$\sum_{n=1}^{\infty}c_{n}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\;+\;d_{0}\;+\;\sum_{n=1}^{\infty}d_{n}\,\Phi_{0}(n\pi x)\;=\;0\quad\text{in }L^{2}(0,1).$ (A.20)
Apply the orthogonal projection $P_{0}$ onto the subspace of functions that are odd with respect to $x=\tfrac{1}{2}$ to both sides. Recalling that $\Phi_{0}(n\pi x)=\sin^{2}(n\pi x)$ is even with respect to $x=\tfrac{1}{2}$, we obtain
$P_{0}\!\left(\sum_{n=1}^{\infty}c_{n}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\right)=0.$ (A.21)
We first show that the family
$\bigl(P_{0}\,\Phi_{2}\bigl((n+1)\pi x\bigr)\bigr)_{n\geq 1}$ (A.22)
is linearly independent. By Lemma 4.3 (ii), this reduces to proving that
$P_{0}\,T_{2}^{*}\!\left(\sum_{n=1}^{\infty}c_{n}\cos\bigl((2n+2)\pi x\bigr)\right)=0\quad\Longrightarrow\quad c_{n}=0\ \text{for all }n.$ (A.23)
Indeed, once $c_{n}=0$ for all $n$, the original relation simplifies to $d_{0}+\sum_{n=1}^{\infty}d_{n}\,\Phi_{0}(n\pi x)=0,$ which forces $d_{n}=0$ for all $n$ (and $d_{0}=0$ as well). This yields the claim.
Define
$f(x)=\sum_{n=1}^{\infty}c_{n}\cos\big((2n+2)\pi x\big).$ (A.24)
The above identity then reads $P_{0}T_{2}^{*}f=0$, that is, both $f$ and $T_{2}^{*}f$ are even with respect to $x=\tfrac{1}{2}$. Moreover, $f$ is orthogonal to the constants and to $\cos(2\pi x)$. By Lemma A.1, it follows that $f\equiv 0$, which proves the result. ∎
Appendix B Proof of Theorem 1.1
@@
$u^{\prime\prime}(r)+\left(\frac{\ell(\ell+1)}{r^{2}}+\hat{q}(r)\right)u(r)=\lambda\,u(r),\qquad r>0,$ (B.1)
@@
For each $\nu=\ell+\tfrac{1}{2}$, we denote by $\varphi(r,\nu,\lambda)$ the regular solution of (B.1), characterized by the boundary condition
$\varphi(r,\nu,\lambda)\sim r^{\nu+\frac{1}{2}},\qquad r\to 0.$ (B.2)
@@
$\varphi_{0}(r,\nu,\lambda)=2^{\nu}\,\lambda^{-\nu/2}\,\Gamma(\nu+1)\,\sqrt{r}\,J_{\nu}(\sqrt{\lambda}\,r),$ (B.3)
@@
The existence and uniqueness of a regular solution are ensured by the following theorem, proved in [12, Lemma 2.2] (see also [13, Lemma B.2]) for locally integrable potentials $q$ on $(0,+\infty)$ under the additional assumption $\int_{0}^{1}r\,|q(r)|\,dr<\infty.$
Theorem B.1.
There exists a unique regular solution $\varphi(r,\nu,\lambda)$ of (B.1) satisfying:
@@
    $|\varphi(r,\nu,\lambda)|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}.$ (B.4)
@@
    $\bigl|\varphi(r,\nu,\lambda)-\varphi_{0}(r,\nu,\lambda)\bigr|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}\int_{0}^{r}\frac{t\,|q(t)|}{1+|\sqrt{\lambda}|\,t}\,dt.$ (B.5)
@@
$J_{\nu}(z)=\frac{(z/2)^{\nu}}{\Gamma(\nu+1)}\prod_{n=1}^{\infty}\left(1-\frac{z^{2}}{j_{\nu,n}^{2}}\right).$ (B.8)
@@
$\varphi_{0}(1,\nu,\lambda)=\prod_{n=1}^{\infty}\left(1-\frac{\lambda}{j_{\nu,n}^{2}}\right).$ (B.9)
@@
$\{\lambda_{\ell,n}(q)\}_{n\geq 1}=\{\tilde{\lambda}_{\ell,n}(\tilde{q})\}_{n\geq 1}.$ (B.10)
@@
$\varphi(1,\nu,\lambda;q)=\varphi(1,\nu,\lambda;\tilde{q}),\qquad\forall\,\lambda\in\mathbb{C}.$ (B.11)
@@
We denote by a tilde all quantities associated with the potential $\tilde{q}$, and, for brevity, we write $\varphi(1,\nu,\lambda;q)$ simply as $\varphi(1,\nu,\lambda)$. Since the Dirichlet spectra coincide, we have $m=\tilde{m}$. By Theorem B.1 (with $r=1$), for real $\lambda=k^{2}$ and $|k|\to\infty$,
$\bigl|\varphi(1,\nu,\lambda)-\varphi_{0}(1,\nu,\lambda)\bigr|=\mathcal{O}(|k|^{-\ell-2}).$ (B.12)
@@
$\varphi_{0}(1,\nu,\lambda)\sim c_{\nu}\,k^{-\ell-1}\cos\!\Big(k-\tfrac{\pi\nu}{2}-\tfrac{\pi}{4}\Big),\qquad k=\sqrt{\lambda}\in\mathbb{R}.$ (B.13)
@@
$k_{n}=\tfrac{\pi}{2}\Big(\nu+\tfrac{1}{2}\Big)+2\pi n,\qquad\lambda_{n}=k_{n}^{2},$ (B.14)
@@
$\varphi(1,\nu,\lambda_{n})\sim\varphi_{0}(1,\nu,\lambda_{n})\sim\tilde{\varphi}(1,\nu,\lambda_{n}).$ (B.15)
@@
One can also define the Jost solutions $f^{\pm}(r,\nu,\lambda)$ as the unique $C^{1}$ solutions with respect to $r$ of (B.1) satisfying the oscillatory asymptotics
$f^{\pm}(r,\nu,\lambda)\sim e^{\pm i\sqrt{\lambda}\,r},\qquad r\to+\infty.$ (B.16)
The proof of this result is given in [13, Lemma B.4] under the assumptions $\int_{0}^{1}r\,|q(r)|\,dr<\infty$ and $\int_{1}^{+\infty}|q(r)|\,dr<\infty.$ In particular, for $\lambda=1$ (which is the case considered in [6] and [20]), and for such potentials supported in $[0,1]$, one has, for all $r\geq 1$ (see [13, Lemma B.3]),
$f^{+}(r,\nu,1)=e^{i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(1)}(r).$ (B.17)
@@
$f^{-}(r,\nu,1)=e^{-i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(2)}(r),$ (B.18)
@@
The pair of Jost solutions $\{f^{+}(r,\nu,1),\,f^{-}(r,\nu,1)\}$ forms a fundamental system of solutions of (B.1). Hence, at the fixed energy $\lambda=1$, the regular solution can be written as
$\varphi(r,\nu,1)=\alpha(\nu)\,f^{+}(r,\nu,1)+\beta(\nu)\,f^{-}(r,\nu,1),$ (B.19)
@@
$\sigma(\nu)=e^{i\pi(\nu+\frac{1}{2})}\,\frac{\alpha(\nu)}{\beta(\nu)}.$ (B.20)
@@
$q=\tilde{q}\quad\text{a.e. on }(0,1).$ (B.22)
@@
Assume now that $q$ and $\tilde{q}$ share the same Dirichlet spectra $\{\lambda_{\ell_{k},n}(q)\}_{k,n\geq 1}$. Since the Jost solutions $f^{\pm}(r,\nu,1)$ never vanish (see [6, Lemma 4.2]) and are independent of the potential at $r=1$, it follows from (B.19) and Lemma B.3 that the corresponding Regge interpolation functions coincide for $\nu_{k}=\ell_{k}+\tfrac{1}{2}$. This concludes the proof of Theorem 1.1.

### v3

@@ 10 Uniqueness result in the case $(\ell_{1},\ell_{2})=(0,3)$
Theoretical Frobenius asymptotics. We recall that the indicial equation at $x=0$ associated with the differential equation in Lemma 10.1 has the following roots:
$\rho=-2,\;1,\;3,\;5,\;6,\;7\,,\qquad\rho=-2\pm 2i\sqrt{11}\,.$
Using Mathematica, we compute the Frobenius expansions associated with each indicial root. Truncating the resulting expressions up to terms $o(1)$ as $x\to 0^{+}$ , we obtain the following local expansion:
[…]
$\;+\;o(1)\,,\qquad x\to 0^{+}\,.$
for some complex constants $A,B,C$ . Note that the remaining Frobenius solutions associated with the real roots $\rho=1,3,5,6,7$ behave like $O(x)$ or higher powers of $x$ , and are therefore absorbed into the $o(1)$ remainder.
Thus, taking either the real or the imaginary part of the previous complex solution $y(x)$ yields a real solution admitting the asymptotic expansion
[…]
$+\frac{1}{x^{2}}\Bigl[\bigl(C_{c}+\widetilde{C}_{c}\,x^{2}\bigr)\cos\!\bigl(2\sqrt{11}\,\ln x\bigr)+\bigl(C_{s}+\widetilde{C}_{s}\,x^{2}\bigr)\sin\!\bigl(2\sqrt{11}\,\ln x\bigr)\Bigr]$
$\;+\;o(1)\,,\qquad x\to 0^{+}\,$,
for some real constants $C_{-2},C_{c},C_{s},\widetilde{C}_{c},\widetilde{C}_{s}$. These constants are not independent: $(C_{c},C_{s})$ uniquely determines $(\widetilde{C}_{c},\widetilde{C}_{s})$ through the linear relation
$\begin{pmatrix}\widetilde{C}_{c}\\[2.84526pt] \widetilde{C}_{s}\end{pmatrix}=\begin{pmatrix}\frac{29}{15}&\frac{13\sqrt{11}}{15}\\[5.69054pt] -\frac{13\sqrt{11}}{15}&\frac{29}{15}\end{pmatrix}\begin{pmatrix}C_{c}\\[2.84526pt] C_{s}\end{pmatrix}$.
We also observe that the oscillatory modes do not contain any term of order $x^{-1}\,$.
Comparison with numerics. Comparing with the numerical ansatz (10.2), we obtain the identifications
$A_{0}=C_{c}\,,\qquad B_{0}=C_{s}\,,\qquad A_{2}=\widetilde{C}_{c}\,,\qquad B_{2}=\widetilde{C}_{s}\,.$
Moreover, the Frobenius structure predicts that
$A_{1}=B_{1}=0\,.$
The fitted numerical values satisfy $|A_{1}|,|B_{1}|\lesssim 10^{-6}$, which is fully compatible with this theoretical prediction. Any nonzero values obtained at lower precision should therefore be interpreted as numerical artefacts.
The numerical fit shows no evidence of a non-oscillatory $x^{-2}$ contribution, which suggests
$C_{-2}=0\,.$
Identifying the leading oscillatory coefficients yields
$C_{c}\approx A_{0}\approx-0.106173\,,\qquad C_{s}\approx B_{0}\approx\phantom{-}0.039403\,.$
Equivalently, writing
$C_{c}\,\cos\!\bigl(2\sqrt{11}\,\ln x\bigr)+C_{s}\,\sin\!\bigl(2\sqrt{11}\,\ln x\bigr)=C_{0}\,\cos\!\bigl(2\sqrt{11}\,\ln x+\phi\bigr)\,$,
we recover
$C_{0}=\sqrt{C_{c}^{2}+C_{s}^{2}}\approx 0.11325\,,\qquad\phi=\arctan\!\Bigl(-\frac{C_{s}}{C_{c}}\Bigr)\approx-2.79\,.$
Finally, the Frobenius expansion implies that the constant-order oscillatory coefficients $\widetilde{C}_{c},\widetilde{C}_{s}$ are not independent. Using the fitted values of $C_{c}$ and $C_{s}$ , this relation predicts
$\widetilde{C}_{c}^{\mathrm{pred}}\approx-0.0917\,,\qquad\widetilde{C}_{s}^{\mathrm{pred}}\approx\phantom{-}0.3819\,.$
These values are in excellent agreement with the numerical coefficients
$\widetilde{C}_{c}\approx A_{2}\approx-0.09236\,,\qquad\widetilde{C}_{s}\approx B_{2}\approx\phantom{-}0.38189\,$,
with relative discrepancies of approximately $0.7\,\%$ for $\widetilde{C}_{c}$ and $3\times 10^{-4}\,\%$ for $\widetilde{C}_{s}$ . This quantitative agreement provides strong numerical evidence for the presence of the quadratic correction in the oscillatory Frobenius modes.
Role of the numerical analysis. At this stage, it is important to clarify the role played by the numerical computations. They are used solely to detect that the solution $y$ of the eighth–order equation is not bounded near $x=0$ and $x=1$ . This observation implies that at least one of the singular Frobenius modes associated with the indicial roots
$\rho=-2\qquad\text{or}\qquad\rho=-2\pm 2i\sqrt{11}$
is present in the expansion of $y$ near the endpoints. No numerical information on the corresponding coefficients is required for this conclusion.
In particular, the proof that the primitive $\zeta$ does not belong to $L^{2}(0,1)$ relies only on the qualitative oscillatory behaviour induced by these Frobenius modes, and not on the explicit values of the constants appearing in the asymptotic expansion. The numerical computation of the coefficients $(C_{c},C_{s},\widetilde{C}_{c},\widetilde{C}_{s})$ should therefore be viewed as an additional quantitative illustration, confirming the Frobenius structure predicted by the theory.
Non–square integrability of the primitive. Let $\zeta$ be a primitive of $y$, namely
$\zeta(x):=\int_{1/2}^{x}y(s)\,ds\,,\qquad x\in(0,1).$
At this stage, and for the sake of generality, we do not assume that $C_{-2}=0$ . Integrating the leading terms in the asymptotic expansion of $y$ as $x\to 0^{+}$ yields
$\zeta(x)=\frac{-120\,C_{-2}}{x}+\frac{C_{0}}{45}\,\frac{-\cos\!\bigl(2\sqrt{11}\,\ln x+\phi\bigr)+2\sqrt{11}\,\sin\!\bigl(2\sqrt{11}\,\ln x+\phi\bigr)}{x}+O(1)\,,\qquad x\to 0^{+}\,.$
Setting $t=-\ln x$ , we can rewrite this behaviour as
$\zeta(e^{-t})=e^{t}\,A(t)+O(1)\,,\qquad t\to+\infty,$
where the function $A$ is bounded, periodic, and explicitly given by
$A(t):=-120\,C_{-2}+\frac{C_{0}}{45}\Bigl(-\cos\!\bigl(2\sqrt{11}\,t-\phi\bigr)-2\sqrt{11}\,\sin\!\bigl(2\sqrt{11}\,t-\phi\bigr)\Bigr)\,.$
Since $A\not\equiv 0$ and $A$ is continuous and periodic, there exist constants $\delta>0$ and $c_{0}>0$ , and a sequence $t_{n}\to+\infty$ , such that
$|A(t)|\geq c_{0}\qquad\text{for all }t\in I_{n}:=[t_{n},t_{n}+\delta]\,.$
Consequently,
$\int_{0}^{1}|\zeta(x)|^{2}\,dx=\int_{0}^{+\infty}|\zeta(e^{-t})|^{2}\,e^{-t}\,dt=\int_{0}^{+\infty}e^{t}\,|A(t)|^{2}\,dt=+\infty\,.$
because
$\int_{0}^{+\infty}e^{t}\,|A(t)|^{2}\,dt\;\geq\;\sum_{n}\int_{I_{n}}e^{t}\,|A(t)|^{2}\,dt\;\geq\;c_{0}^{2}\sum_{n}\int_{t_{n}}^{t_{n}+\delta}e^{t}\,dt=+\infty\,.$
This proves that $\zeta\notin L^{2}(0,1)$ .
Step 4: Conclusion. We therefore conclude that one must take $y(\tfrac{1}{2})=0$ . By the triangular system $(S)$, all derivatives of $y$ at $x=\tfrac{1}{2}$ up to order $7$ vanish. The Cauchy–Lipschitz theorem then implies that $y\equiv 0$ on $(0,1)$. Consequently, $\zeta\equiv 0$ on $(0,1)$ .
In particular, this implies that the differential of the spectral map associated with the pair $(\ell_{1},\ell_{2})=(0,3)$ is injective .
11 Conclusion
The techniques developed in this paper rely on the Kneser–Sommerfeld formula. In the case $(\ell_{1},\ell_{2})=(0,3)\,$, the analysis required the use of computer-assisted numerical computations. Even with the assistance of computer algebra, the study of the resulting differential equations already becomes rapidly cumbersome in the case $(\ell_{1},\ell_{2})=(1,3)\,$. Larger values of $(\ell_{1},\ell_{2})$ currently seem out of reach, which suggests that a more conceptual approach is needed.
Appendix A Proof of Theorem 1.1
@@ 11 Conclusion
$u^{\prime\prime}(r)+\left(\frac{\ell(\ell+1)}{r^{2}}+\hat{q}(r)\right)u(r)=\lambda\,u(r)\,,\qquad r>0,$ (A.1)
@@
For each $\nu=\ell+\tfrac{1}{2}$, we denote by $\varphi(r,\nu,\lambda)$ the regular solution of (A.1), characterized by the boundary condition
$\varphi(r,\nu,\lambda)\sim r^{\nu+\frac{1}{2}}\,,\qquad r\to 0\,.$ (A.2)
@@
$\varphi_{0}(r,\nu,\lambda)=2^{\nu}\,\lambda^{-\nu/2}\,\Gamma(\nu+1)\,\sqrt{r}\,J_{\nu}(\sqrt{\lambda}\,r)\,$, (A.3)
@@
The existence and uniqueness of a regular solution are ensured by the following theorem, proved in [15, Lemma 2.2] (see also [16, Lemma B.2]) for locally integrable potentials $q$ on $(0,+\infty)$ under the additional assumption $\int_{0}^{1}r\,|q(r)|\,dr<\infty\,.$
Theorem A.1.
There exists a unique regular solution $\varphi(r,\nu,\lambda)$ of (A.1) satisfying:
@@
    $|\varphi(r,\nu,\lambda)|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}\,.$ (A.4)
@@
    $\bigl|\varphi(r,\nu,\lambda)-\varphi_{0}(r,\nu,\lambda)\bigr|\;\leq\;C\left(\frac{r}{1+|\sqrt{\lambda}|\,r}\right)^{\ell+1}e^{|\Im\sqrt{\lambda}|\,r}\int_{0}^{r}\frac{t\,|q(t)|}{1+|\sqrt{\lambda}|\,t}\,dt\,.$ (A.5)
@@
$J_{\nu}(z)=\frac{(z/2)^{\nu}}{\Gamma(\nu+1)}\prod_{n=1}^{\infty}\left(1-\frac{z^{2}}{j_{\nu,n}^{2}}\right)\,.$ (A.8)
@@
$\varphi_{0}(1,\nu,\lambda)=\prod_{n=1}^{\infty}\left(1-\frac{\lambda}{j_{\nu,n}^{2}}\right)\,.$ (A.9)
@@
$\{\lambda_{\ell,n}(q)\}_{n\geq 1}=\{\lambda_{\ell,n}(\tilde{q})\}_{n\geq 1}\,.$ (A.10)
@@
$\varphi(1,\nu,\lambda;q)=\varphi(1,\nu,\lambda;\tilde{q})\,,\qquad\forall\,\lambda\in\mathbb{C}\,.$ (A.11)
@@
We denote by a tilde all quantities associated with the potential $\tilde{q}$, and, for brevity, we write $\varphi(1,\nu,\lambda;q)$ simply as $\varphi(1,\nu,\lambda)$. Since the Dirichlet spectra coincide, we have $m=\tilde{m}$ . By Theorem A.1 (with $r=1$), for real $\lambda=k^{2}$ and $|k|\to\infty$ ,
$\bigl|\varphi(1,\nu,\lambda)-\varphi_{0}(1,\nu,\lambda)\bigr|=\mathcal{O}(|k|^{-\ell-2})\,.$ (A.12)
@@
$\varphi_{0}(1,\nu,\lambda)\sim c_{\nu}\,k^{-\ell-1}\cos\!\Big(k-\tfrac{\pi\nu}{2}-\tfrac{\pi}{4}\Big)\,,\qquad k=\sqrt{\lambda}\in\mathbb{R}\,.$ (A.13)
@@
$k_{n}=\tfrac{\pi}{2}\Big(\nu+\tfrac{1}{2}\Big)+2\pi n\,,\qquad\lambda_{n}=k_{n}^{2}\,$, (A.14)
@@
$\varphi(1,\nu,\lambda_{n})\sim\varphi_{0}(1,\nu,\lambda_{n})\sim\tilde{\varphi}(1,\nu,\lambda_{n})\,.$ (A.15)
@@
One can also define the Jost solutions $f^{\pm}(r,\nu,\lambda)$ as the unique $C^{1}$ solutions with respect to $r$ of (A.1) satisfying the oscillatory asymptotics
$f^{\pm}(r,\nu,\lambda)\sim e^{\pm i\sqrt{\lambda}\,r}\,,\qquad r\to+\infty\,.$ (A.16)
The proof of this result is given in [16, Lemma B.4] under the assumptions $\int_{0}^{1}r\,|q(r)|\,dr<\infty$ and $\int_{1}^{+\infty}|q(r)|\,dr<\infty\,.$ In particular, for $\lambda=1$ (which is the case considered in [7] and [24]), and for such potentials supported in $[0,1]$, one has, for all $r\geq 1$ (see [16, Lemma B.3]),
$f^{+}(r,\nu,1)=e^{i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(1)}(r)$ (A.17)
@@
$f^{-}(r,\nu,1)=e^{-i(\nu+\frac{1}{2})\frac{\pi}{2}}\,\sqrt{\frac{\pi r}{2}}\,H_{\nu}^{(2)}(r)\,$, (A.18)
@@
The pair of Jost solutions $\{f^{+}(r,\nu,1),\,f^{-}(r,\nu,1)\}$ forms a fundamental system of solutions of (A.1). Hence, at the fixed energy $\lambda=1$, the regular solution can be written as
$\varphi(r,\nu,1)=\alpha(\nu)\,f^{+}(r,\nu,1)+\beta(\nu)\,f^{-}(r,\nu,1)\,$, (A.19)
@@
$\sigma(\nu)=e^{i\pi(\nu+\frac{1}{2})}\,\frac{\alpha(\nu)}{\beta(\nu)}\,.$ (A.20)
@@
$q=\tilde{q}\quad\text{a.e. on }(0,1)\,.$ (A.22)
@@
Assume now that $q$ and $\tilde{q}$ share the same Dirichlet spectra $\{\lambda_{\ell_{k},n}(q)\}_{k,n\geq 1}$ . Since the Jost solutions $f^{\pm}(r,\nu,1)$ never vanish (see [7, Lemma 4.2]) and are independent of the potential at $r=1$ , it follows from (A.19) and Lemma A.3 that the corresponding Regge interpolation functions coincide for $\nu_{k}=\ell_{k}+\tfrac{1}{2}$ . This concludes the proof of Theorem 1.1.
Appendix B Proof of Theorem 2.5
In this section, we prove that the range of the differential of the spectral map at the zero potential is closed in the cases $(\ell_{1},\ell_{2})=(0,1)\,,\ (1,2)$ and $(0,3)$ . Moreover, in the case $(0,1)$, we give an explicit description of this range.
We first recall that in both cases $(0,1)$ , $(1,2)$ and $(0,3)$ , the differential of the spectral map at $q=0$ is injective. Equivalently, the intersection of the tangent spaces to the corresponding isospectral manifolds at $q=0$ is trivial. Following the terminology of Shubin Christ [29] and Carlson–Shubin [5], we introduce
$W\;:=\;T_{0}\mathcal{M}_{\ell_{1}}\;\cap\;T_{0}\mathcal{M}_{\ell_{2}}\,$,
where $\mathcal{M}_{\ell}$ denotes the isospectral set associated with the Dirichlet spectrum for angular momentum $\ell$ . Injectivity of the differential is precisely the statement that
$W=\{0\}\,.$
Thus, in the case $(\ell_{1},\ell_{2})=(0,1)$ , the first part of Theorem 2.5 follows from ([29], Lemma 5), while for $(\ell_{1},\ell_{2})=(1,2)$ or $(0,3)$, it follows from the analysis carried out in the proof of ([5], Theorem 5.6).
We have thus shown that in the cases $(\ell_{1},\ell_{2})=(0,1)\,$, $(\ell_{1},\ell_{2})=(1,2)\,$ and $(\ell_{1},\ell_{2})=(0,3)$ , the differential of the spectral map at $q=0$ is injective and has closed range. In other words, the operator $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ is semi-Fredholm (see [12, p. 230]). Moreover, if an operator $B$ is Fredholm or semi-Fredholm, then any compact perturbation remains Fredholm or semi-Fredholm and the Fredholm index is preserved (see [12, Theorem 5.26]).
We now use these theoretical results to refine the analysis in the case $(0,1)$. To this end, we set
$\Phi_{\ell}(x):=\frac{\pi}{2}\,x\,J_{\ell+\frac{1}{2}}(x)^{2}=(j_{\ell}(x))^{2}\,.$
First, we decompose the differential $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}$ into two parts,
$d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}=A_{\ell_{1},\ell_{2}}+K_{\ell_{1},\ell_{2}}\,$,
where we define
$A_{\ell_{1},\ell_{2}}(\zeta)=\Bigg(\langle\zeta,1\rangle,\;$ $\bigl(\langle\zeta,\;2\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi\,\cdot\right)-1\rangle\bigr)_{n\geq 1}\,$,
$\bigl(\langle\zeta,\;2\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi\,\cdot\right)-1\rangle\bigr)_{n\geq 1}\Bigg)\,$, (B.1)
and
$K_{\ell_{1},\ell_{2}}(\zeta)=\Bigg(0,\;$ $\bigl(\langle\zeta,\;g_{\ell_{1},n}^{2}-2\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi\,\cdot\right)\rangle\bigr)_{n\geq 1}\,$,
$\bigl(\langle\zeta,\;g_{\ell_{2},n}^{2}-2\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi\,\cdot\right)\rangle\bigr)_{n\geq 1}\Bigg)\,.$ (B.2)
According to Theorem 2.1 and Corollary 2.2 in Serier [28], one has
$\|g_{\ell,n}^{2}-2\Phi_{\ell}(j_{\nu,n}\,\cdot)\|_{L^{2}(0,1)}=\mathcal{O}\!\left(\frac{1}{n}\right)\,.$
Since $|\Phi^{\prime}_{\ell}(x)|\leq C_{\ell}$ uniformly on $\mathbb{R}$ and
[…]
it follows that
$\sum_{n\geq 1}\bigl\|g_{\ell,n}^{2}-2\Phi_{\ell}\bigl((n+\tfrac{\ell}{2})\pi\,\cdot\bigr)\bigr\|_{L^{2}(0,1)}^{2}<\infty\,.$
Therefore, by the Cauchy–Schwarz inequality, the operator $K_{\ell_{1},\ell_{2}}$ is Hilbert-Schmidt, and hence compact, from $L^{2}(0,1)$ into $\mathbb{R}\times\ell^{2}(\mathbb{N})\times\ell^{2}(\mathbb{N})$ . As a consequence of the stability of semi-Fredholm operators under compact perturbations, it follows that, for $(\ell_{1},\ell_{2})=(0,1)$, $(\ell_{1},\ell_{2})=(1,2)\,$ or $(0,3)$ the range of $A_{\ell_{1},\ell_{2}}\,$, which is a compact perturbation of $d_{0}\mathcal{S}_{\ell_{1},\ell_{2}}\,$, is also closed.
Now, according to [27, Corollary 5.2], in the case $(\ell_{1},\ell_{2})=(0,1)\,$, the family
$\left(1,\;\bigl(\Phi_{\ell_{1}}((n+\tfrac{\ell_{1}}{2})\pi\,\cdot)\bigr)_{n\geq 1},\;\bigl(\Phi_{\ell_{2}}((n+\tfrac{\ell_{2}}{2})\pi\,\cdot)\bigr)_{n\geq 1}\right)$
is referred to as a basis of $L^{2}(0,1)$ [^3] This statement should be interpreted with care. In the sense of Gohberg-Krein [10], this only means that the above family is complete and $\omega$-linearly independent in $L^{2}(0,1)$ This property does not, in general, imply that this family is a Krein basis or a Riesz basis.. We define
$\mathcal{F}_{\ell_{1},\ell_{2}}=\Bigl(f_{0},\ (f_{1,n})_{n\geq 1}\,,\ (f_{2,n})_{n\geq 1}\Bigr)\subset L^{2}(0,1)\,$,
where
$f_{0}(r)=1,\qquad f_{1,n}(r)=2\Phi_{\ell_{1}}\!\left(\bigl(n+\tfrac{\ell_{1}}{2}\bigr)\pi r\right)-1\,,\qquad f_{2,n}(r)=2\Phi_{\ell_{2}}\!\left(\bigl(n+\tfrac{\ell_{2}}{2}\bigr)\pi r\right)-1\,.$
The family $\mathcal{F}_{\ell_{1},\ell_{2}}$ remains complete and algebraically linearly independent in $L^{2}(0,1)$ . With this notation, the operator $A_{\ell_{1},\ell_{2}}$ can be written as
$A_{\ell_{1},\ell_{2}}(\zeta)=\Bigl(\langle\zeta,f_{0}\rangle,\ (\langle\zeta,f_{1,n}\rangle)_{n\geq 1},\ (\langle\zeta,f_{2,n}\rangle)_{n\geq 1}\Bigr)\,$,
and is therefore injective. Recall that $c_{00}$ denotes the space of finitely supported sequences. Since
$\mathbb{R}\times c_{00}\times c_{00}\quad\text{is dense in}\quad\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\,$,
it follows from a straightforward algebraic argument that the range of $A_{\ell_{1},\ell_{2}}$ is also dense in $\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})$. Since this range is closed, as shown above, we conclude that
$\operatorname{Ran}A_{\ell_{1},\ell_{2}}=\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\,.$
In particular, the operator $A_{\ell_{1},\ell_{2}}$ is Fredholm of index zero.
Finally, since $d_{0}\mathcal{S}_{0,1}$ is a compact perturbation of $A_{0,1}$, it follows that $d_{0}\mathcal{S}_{0,1}$ is also Fredholm with the same index. Because it is injective, it must therefore be an isomorphism from $L^{2}(0,1)$ onto $\mathbb{R}\times\ell^{2}_{\mathbb{R}}(\mathbb{N})\times\ell^{2}_{\mathbb{R}}(\mathbb{N})$ . This completes the proof of Theorem 2.5.